OpenStudy (anonymous):

whats the solution: 3/2 - 1/x-4 = 1/2x-8 2 3 4 5

OpenStudy (radar):

5

OpenStudy (radar):

3/2 -1/(x-4) do the subtraction getting (3x-14)/(2x-8) =1/(2x-8) cross multiply 2x-8=(3x-14)(2x-8) Divide both sides by (2x-8) 1=3x-14 3x=15 x=5